# Thermodynamic inference in molecular motors: a Martingale approach

### v1

@@ III.2 Inference results from the IFT at stopping times with approximations
It is worth remarking that in all the above tests, we used a small number of sampled trajectories ($10^{2}$) for the inferred values on purpose. The aim of this work is to provide reliable and feasible thermodynamic inference strategies. Thus, obtaining accurate results cannot rely on sampling a potentially unfeasible large number of trajectories for experimentalists. In this context, it is important to keep in mind the time that these measurements could eventually imply in a realistic experimental setup. In the F1-ATPase case, the motor completes a cycle in the forward (anticlockwise) hydrolysis direction at a time of the order of $0.1\leavevmode\nobreak\ \mathrm{s}$. The most time-consuming situation we considered corresponds to an upper threshold with $N_{+}=30$, which requires, on average, $3\leavevmode\nobreak\ \mathrm{s}$ to complete the stopping protocol. Recording the motion of $10^{2}$ trajectories would then take $300\leavevmode\nobreak\ \mathrm{s}=5\leavevmode\nobreak\ \mathrm{min}$. That means that the proposed inferred method would be feasible under experimental conditions, since measurements can be obtained rather fast. To prove the reliability of the results with only $10^{2}$ trajectories, we show in Appendix D the dependence of our results with the number of trajectories used.
@@ III.3 Bounds to maximal excursions of entropy production
In this case, recording the motion of $3\cdot 10^{3}$ trajectories, stopped at $\tau=0.5\leavevmode\nobreak\ \mathrm{s}$, would take $25\leavevmode\nobreak\ \mathrm{min}$, which is a rather short time for an experimental measurement. Again, the experimental proposal requires simple and fast measurements. The dependence of the reported results with the number of simulated trajectories used for inference is commented in Appendix D. Similarly to the previous results for the IFT at stopping times in Secs. III.1 and III.2, the inferred values and the accuracy of the method do not change significantly when increasing the number of trajectories by a factor $10$, to $10^{3}$.
As for the strategies that use the IFT at stopping times, the above analysis could have been shaped for the case in which the motor is biased to rotate on average in the synthesis direction (clockwise). The corresponding expression useful for inference can be obtained from Eq. (16) by exchanging $W_{\mathrm{ext}}$ and $\Delta\mu$, while the macroscopic efficiency would be in this case $\eta_{\mathrm{syn}}=\Delta\mu/W_{\mathrm{ext}}$. More details about this parallel situation are given in Appendix C.
@@ IV Conclusions
The strategy is based on a stopping protocol that allows us to know exactly the entropy production of the stopped trajectories by starting to record the motor motion after a first jump in the signalling transition [see Fig. 1]. A rigorous application of such strategy requires to take into account the presence of absolute irreversibility, which may be often difficult to obtain. Our simulations of the inference strategy for the rotational work performed by the F1-ATPase motor give results compatible with the true value within one standard deviation and with a relative deviation below $0.06\%$. Nevertheless, we also showed that working directly in the NESS, and neglecting both absolute irreversibility and an initial transient, the strategy is simplified while still giving accurate results. Using this approximate method, we obtained an inferred value that is also compatible within one standard deviation with the true value and with a relative deviation below $0.12\%$.
@@
As a last remark, we highlight that, while we have focused on the illustrative example of the F1-ATPase rotatory motor, the strategies introduced in this paper are general and can be applied to other molecular motors acting on cycles with at least one signaling transition. In particular, it would be interesting to apply these strategies for the case of myosin (see, for instance, the model proposed for its dynamics in Ref. [58]), for discrete-state models of kinesin (see e.g. Refs. [59, 60, 61]), or to other biochemical processes described as cycles [62, 63, 64, 65], which may include information flows [66, 67]. Moreover, our inference methods may be applied beyond the case of biomolecular machines, such as for synthetic thermal machines operating in steady-state conditions implemented in nanoelectronic devices [68, 69, 70, 71, 72, 73]. These methods might also be extended to the quantum case by using extensions of the martingale theory for stochastic entropy production for monitored quantum systems [74, 75], with possible applications in quantum thermal machines [76, 77, 78, 79].
@@ Acknowledgments
ANR was funded by the Consejo Superior de Investigaciones Científicas (CSIC) JAE Intro 2023 program. GM acknowledges support from the Ramón y Cajal program RYC2021-031121-I funded by MICIU/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR and the María de Maeztu project CEX2021-001164-M for Units of Excellence, funded by MICIU/AEI/10.13039/501100011033/FEDER, UE.
Appendix A Master equation and steady state
$\begin{gathered}\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}\rho_{x_{0}}(t)\\ \rho_{x_{1}}(t)\\ \rho_{x_{2}}(t)\end{pmatrix}=\mathbb{K}\begin{pmatrix}\rho_{x_{0}}(t)\\ \rho_{x_{1}}(t)\\ \rho_{x_{2}}(t)\end{pmatrix},\end{gathered}$ (23)
$\begin{split}N\pi_{x_{0}}=K_{x_{2}x_{0}}K_{x_{1}x_{2}}+K_{x_{1}x_{0}}(K_{x_{2}x_{0}}+K_{x_{2}x_{1}}),\\ N\pi_{x_{1}}=K_{x_{2}x_{0}}K_{x_{0}x_{1}}+K_{x_{2}x_{1}}(K_{x_{0}x_{1}}+K_{x_{0}x_{2}}),\\ N\pi_{x_{2}}=K_{x_{1}x_{0}}K_{x_{0}x_{2}}+K_{x_{1}x_{2}}(K_{x_{0}x_{1}}+K_{x_{0}x_{2}}),\\ \end{split}$ (25)
We remark that the stationary probabilities $\{\pi_{x_{0}},\pi_{x_{1}},\pi_{x_{2}}\}$ calculated above are often not accessible from the observation of the hydrolysis and synthesis transition. However, in Appendix B we show a method to adapt the information from the observable transitions to infer information of the other transitions, obtaining a reliable description of the motor dynamics. Those are also the results we used alongside the theoretical rates in Eq. (19) and the experimental values provided in Tab. 1 to simulate the motor dynamics.
We use the expression above in Appendix B to set the time scale of the motor dynamics in the absence of external torque, leading to $W_{\mathrm{ext}}=0$. (We notice that in all figures we instead have $W_{\mathrm{ext}}=17\leavevmode\nobreak\ k_{\mathrm{B}}T$.)
Appendix B Experimental values for the parameters
Finally, the only available measurement of the time scale of the motor dynamics is its mean rotational speed. However, a full 360º rotation includes three net hydrolysis or synthesis steps in our model, which implies three cycles in either direction once. Given the lack of information about the rates of individual transitions, the first approximation was to consider $k_{\mathrm{ATP}}=k_{\mathrm{ADP}}=k_{HS}=\gamma$, with $\gamma$ the parameter that establishes the time scale of our simulations. Under normal conditions, the motor is expected to work in its stationary state, so we need the rate at which cycles are completed in the stationary regime and relate it to the experimental measurements. We obtained in Appendix A that the rate of cycles is given by Eq. (26). According to the first experimental characterization of the F1-ATPase in Ref. [18], in the absence of external forces ($W_{\mathrm{ext}}=0$) the motor rotates in the hydrolysis direction doing around $K_{\rm revs}=17$ r.p.s. We hence approximate the value of the rate $\gamma$ as:
$K_{x_{1}x_{2}}\pi_{x_{1}}-K_{x_{2}x_{1}}\pi_{x_{2}}=$
Appendix C Bounds for the output mechanical work
@@
[…]
Note that the macroscopic efficiency is time-independent. We can then write $W^{\mathrm{c}}(t)\right>=\eta^{-1}\left<W^{\mathrm{m}}(t)\right>$, and express the chemical work in a trajectory up to a time $t$ by:
@@
being $\xi(t)$ a non-extensive in time stochastic variable with zero mean, $\xi(t)\right>=0$. Going back to the total entropy production, using Eq. (31), we can now rewrite it as:
@@
$\begin{split}&P\left[W^{\rm m}_{\mathrm{max}}(t)\leq w_{+}\right]=1-P\left[W^{\rm m}_{\max}(t)\geq w_{+}\right]\\ &=1-P\left[\mathrm{max}_{t\leq\tau}S_{\mathrm{tot}}(t)\geq s_{+}\right]=1-\alpha,\\ \end{split}$ (33)
@@
and where we denoted $W^{\rm m}_{\mathrm{max}}(t)=\max_{t\leq\tau}W^{\rm m}(t)$ the maximum of $W(t)$ within the interval $[0,\tau]$. Equivalently, for the lower bound we have:
@@
and where we analogously denoted $W_{\min}^{\mathrm{m}}(t)=\mathrm{min}_{t\leq\tau}W^{\mathrm{m}}(t)$ the minimum of the mechanical work during the interval $[0,\tau]$.
@@
which, after some basic algebra, give the following result for the optimal famility of thresholds not surpassed during the interval $[0,\tau]$ with probability $1-\alpha$:
@@
Notice that the optimal upper bound still needs a reconstruction of the mean of the exponential of the entropy production through $\lambda(1)$. However, the optimal lower bound is achieved for $p=-1$, and we have $\lambda(-1)=1$ from the standard IFT [81]. Therefore, in this case the bound saturates to a time-independent value:

### v2

@@ III.2 Inference results from the IFT at stopping times with approximations
It is worth remarking that in all the above tests, we used a small number of sampled trajectories ($10^{2}$) for the inferred values on purpose. The aim of this work is to provide reliable and feasible thermodynamic inference strategies. Thus, obtaining accurate results cannot rely on sampling a potentially unfeasible large number of trajectories for experimentalists. In this context, it is important to keep in mind the time that these measurements could eventually imply in a realistic experimental setup. In the F1-ATPase case, the motor completes a cycle in the forward (anticlockwise) hydrolysis direction at a time of the order of $0.1\leavevmode\nobreak\ \mathrm{s}$. The most time-consuming situation we considered corresponds to an upper threshold with $N_{+}=30$, which requires, on average, $3\leavevmode\nobreak\ \mathrm{s}$ to complete the stopping protocol. Recording the motion of $10^{2}$ trajectories would then take $300\leavevmode\nobreak\ \mathrm{s}=5\leavevmode\nobreak\ \mathrm{min}$. That means that the proposed inferred method would be feasible under experimental conditions, since measurements can be obtained rather fast. To prove the reliability of the results with only $10^{2}$ trajectories, we show in Appendix E the dependence of our results with the number of trajectories used.
@@ III.3 Bounds to maximal excursions of entropy production
In this case, recording the motion of $3\cdot 10^{3}$ trajectories, stopped at $\tau=0.5\leavevmode\nobreak\ \mathrm{s}$, would take $25\leavevmode\nobreak\ \mathrm{min}$, which is a rather short time for an experimental measurement. Again, the experimental proposal requires simple and fast measurements. The dependence of the reported results with the number of simulated trajectories used for inference is commented in Appendix E. Similarly to the previous results for the IFT at stopping times in Secs. III.1 and III.2, the inferred values and the accuracy of the method do not change significantly when increasing the number of trajectories by a factor $10$, to $10^{3}$.
As for the strategies that use the IFT at stopping times, the above analysis could have been shaped for the case in which the motor is biased to rotate on average in the synthesis direction (clockwise). The corresponding expression useful for inference can be obtained from Eq. (16) by exchanging $W_{\mathrm{ext}}$ and $\Delta\mu$, while the macroscopic efficiency would be in this case $\eta_{\mathrm{syn}}=\Delta\mu/W_{\mathrm{ext}}$. More details about this parallel situation are given in Appendix B.
@@ IV Conclusions
The strategy is based on a stopping protocol that allows us to know exactly the entropy production of the stopped trajectories by starting to record the motor motion after a first jump in the signalling transition [see Fig. 1]. A rigorous application of such strategy requires to take into account the presence of absolute irreversibility, which may be often difficult to obtain. Our simulations of the inference strategy for the rotational work performed by the F1-ATPase motor give results compatible with the true value within one standard deviation and with a relative deviation below $0.06\%$. Nevertheless, we also showed that working directly in the NESS, and neglecting both absolute irreversibility and an initial transient, the strategy is simplified while still giving accurate results. Using this approximate method, we obtained an inferred value that is also compatible within one standard deviation with the true value and with a relative deviation below $0.12\%$. Importantly, finite statistics estimation of the IFT at stopping times may converge to its theoretical value much faster than the conventional IFT [50]. Intuitively, this can be made by setting the thresholds to avoid rare trajectories with e.g. large negative entropy production. As a consequence, this strategy becomes suitable for situations with a limited sample of trajectories.
@@
The above numbers for sampling times are comparable to the ones obtained in the experiments using the detailed fluctuation theorem for inference purposes as reported in Ref. [27], while the strategies presented here allow more flexibility in the choice of quantities to be inferred and seem to provide relative errors up to 5 times smaller. In this context, it would be also interesting to explore inference methods based on recently derived detailed fluctuation theorems applicable at stopping times [61].
As a last remark, we highlight that, while we have focused on the illustrative example of the F1-ATPase rotatory motor, the strategies introduced in this paper are general and can be applied to other molecular motors acting on cycles with at least one signaling transition. In particular, it would be interesting to apply these strategies for the case of myosin (see, for instance, the model proposed for its dynamics in Ref. [62]), for discrete-state models of kinesin (see e.g. Refs. [63, 64, 65]), or to other biochemical processes described as cycles [66, 67, 68, 69], which may include information flows [70, 71, 72, 73]. Moreover, our inference methods may be applied beyond the case of biomolecular machines, such as for synthetic thermal machines operating in steady-state conditions implemented in nanoelectronic devices [74, 75, 76, 77, 78, 79]. These methods might also be extended to the quantum case by using extensions of the martingale theory for stochastic entropy production for monitored quantum systems [80, 81], with possible applications in quantum thermal machines [82, 83, 84, 85].
@@ Acknowledgments
We wish to thank Léa Bresque, Debraj Das and Édgar Roldán for insighful and fruitful discussions. ANR was funded by the Consejo Superior de Investigaciones Científicas (CSIC) JAE Intro 2023 program. GM acknowledges support from the Ramón y Cajal program RYC2021-031121-I funded by MICIU/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR, the CoQuSy project PID2022-140506NB-C21 and 22, and the María de Maeztu project CEX2021-001164-M for Units of Excellence, funded by MICIU/AEI/10.13039/501100011033/FEDER, UE.
The supporting data and codes for this article are openly available from the Ref. [86].
Appendix A Martingale theory in stochastic thermodynamics
Martingales are a class of stochastic processes characterized by the property that their expected value at any time, $\tau$, equals its value at a previous time, $t$, when the expectation is conditioned on observations up to that previous time $t$. That is, the process $M(\tau)$ is a martingale if it verifies:
$\left<M(\tau)|X_{[0,t]}\right>=M(t),$ (23)
with $t\leq\tau$. As a consequence, martingales are unbiased processes that maintain a constant mean over time. Two closely related definitions are submartingales, for which the value of the process at time $t$ is a lower bound of the conditional expectation at $\tau$, $\left<M(\tau)|X_{[0,t]}\right>\geq M(t)$, and supermartingales, for which it is an upper bound, i.e. $\left<M(\tau)|X_{[0,t]}\right>\leq M(t)$.
The introduction of martingales became a breakthrough in probability theory [55], but they have also played an important role in fields like game theory and quantitative finance. In the last years, the use of martingales in the context of stochastic thermodynamics has begun to be explored [45], mainly motivated by the fact that the exponential of (minus) the total entropy production is a martingale in nonequilibrium stationary processes:
$\left<e^{-S_{\mathrm{tot}}(X_{[0,\tau]})}|X_{[0,t]}\right>=e^{-S_{\mathrm{tot}}(X_{[0,t]})}.$ (24)
Indeed, the martingale property above naturally implies the conventional IFT for entropy production at fixed times [14]. This can be seen by applying the martingale property to the onset of the trajectory, when no entropy could have been produced yet, $S_{\mathrm{tot}}(X_{[0,0]})=0$, that is:
$\left<e^{-S_{\mathrm{tot}}(X_{[0,\tau]})}\right$> $=\sum_{x_{0}}\left<e^{-S_{\mathrm{tot}}(X_{[0,\tau]})}|x_{0}\right>\rho_{x_{0}}(0)$ (25)
$=\sum_{x_{0}}e^{-S_{\mathrm{tot}}(X_{[0,0]})}\rho_{x_{0}}(0)=1,$
which implies that the martingale is a stronger property of entropy production than the IFT.
Importantly, all results obtained to characterize martingale processes - see, for instance, Doob’s seminal work [87] - can then be applied to unveil new properties of the statistics of entropy production, which can also be used to further characterize the statistics of other thermodynamic magnitudes of interest. In this context, one useful result of martingales is Doob’s optional stopping theorem. It states that the expected value of a martingale at a (bounded) stopping time, $\mathcal{T}$, is equal to its value at the onset of the trajectory, i.e. $\left<M(\mathcal{T})\right>_{\mathcal{T}}=\left<M(0)\right$>, where the subscript $\mathcal{T}$ in the bracket emphasizes that the (stopped) trajectories in the average may have different lengths.
Applying Doob’s optional stopping theorem to the martingale $e^{-S_{\rm tot}(X_{0,\tau})}$, allows to obtain the IFT at stopping times for stationary processes, which reads [42, 50]:
$\left<e^{-S_{\mathrm{tot}}(X_{[0,\mathcal{T}]})}\right>_{\mathcal{T}}=1,$ (26)
which is valid in a NESS for generic stopping conditions for which $\mathcal{T}<\infty$.
The above results can be generalized to arbitrary out-of-equilibrium processes [46] by noticing the validity of the more general martingale process:
$\left<e^{-S_{\mathrm{tot}}(X_{[0,\tau]})-\delta_{\tau}(\tau)}|X_{[0,t]}\right>=e^{-S_{\mathrm{tot}}(X_{[0,t]})-\delta_{\tau}(t)},$ (27)
with the extra term $\delta_{\tau}(t)$ given in Eq. (2), which vanishes for $t=\tau$ and for stationary dynamics. Applying Doob’s optional stopping theorem to the martingale $e^{-S_{\mathrm{tot}}(X_{[0,t]})-\delta_{\tau}(t)}$ gives us the following IFT at stopping times:
$\left<e^{-S_{\mathrm{tot}}(X_{[0,\mathcal{T}]}-\delta_{\tau}(\mathcal{T}))}\right>_{\mathcal{T}}=1,$ (28)
which is verified in arbitrary non-equilibrium situations and again for generic stopping conditions for which $\mathcal{T}<\infty$. Moreover, in the presence of absolute irreversibility [54], it has been recently shown that Eq. (27) above, becomes a supermartingale [47], and the IFT at stopping times with absolute irreversibily becomes:
$\left<e^{-\left[S_{\mathrm{tot}}(\mathcal{T})-\delta(\mathcal{T})\right]}\right>_{\mathcal{T}}=1-\Gamma$ (29)
with the abolute irreversibility term $0\leq\Gamma\leq 1$ is given in Eq. (4).
Another important universal result for martingales is Doob’s maximal inequality [87], which can be used to bound maximal excursions of entropy production in steady state processes. Besides, once the entropy production is bounded, one can also derive bounds for all thermodynamic quantities that are proportional to it.
More precisely, the fact that $e^{-S_{\mathrm{tot}}}$ is a martingale implies that $e^{qS_{\mathrm{tot}}}$ is a submartingale for any real number $q$ verifying $|q|\geq 1$, that is:
$\left<e^{qS_{\mathrm{tot}}(X_{[0,t]})}|X_{[0,\tau]}\right>\geq e^{qS_{\mathrm{tot}}(X_{[0,\tau]})}.$ (30)
This result can be obtained from Jensen’s inequality for conditional expectations, as follows from Ref. [48]. Let $M(t)$ be a martingale process. Given that $f(x)=x^{-q}$ is a convex function for $|q|\geq 1$, Jensen’s inequality implies that:
$\left<f[M(t)]|X_{[0,\tau]}\right>\geq f\left[\left<M(t)|X_{[0,\tau]}\right>\right]=f[M(\tau)],$ (31)
what proofs the above inequality if identifying the martingale process as the exponential of minus entropy production, $M(t)=e^{-S_{\mathrm{tot}}(X_{[0,t]})}$.
Doob’s maximal inequality states that the probability that the maximum of a non-negative submartingale surpasses a certain value up to a given time is bounded by the mean value of the submartingale at that time:
$P\left(\mathrm{max}_{t\leq\tau}M(t)\geq\lambda\right)\leq\frac{1}{\lambda}\left<M(\tau)\right$>, (32)
with $\lambda>0$. We have already proven that $e^{qS_{\mathrm{tot}}}$, with $|q|\geq 1$ is a submartingale, and, since it is an exponential, we also know that it is non-negative. Using Doob’s maximal inequality we can then bound the probability for a maximal excursion of the above exponentials of entropy production, and then of the entropy production itself.
To do so, it is convenient to work with two subcases of the above exponentials, namely, $e^{\pm pS_{\mathrm{tot}}}$ with $p\geq 1$. Applying Doob’s maximal to these two separate cases:
[…]
where we simplified the notation by writing $S_{\mathrm{tot}}(X_{[0,t]})$ as $S_{\mathrm{tot}}(t)$. Since the exponential $e^{\pm S_{\mathrm{tot}}(t)}$ is an injective, monotonically increasing (+p+p) or decreasing (p-p) function of $S_{\mathrm{tot}}(t)$, there is a one to one correspondence between the value taken by the exponential and the one taken by the entropy production. We can then write $\lambda=e^{\pm ps}$, with $s$ a threshold for $S_{\mathrm{tot}}$. In the case $e^{+pS_{\mathrm{tot}}}$, that functional increases as the entropy production increases, and the maximal excursion of the exponential implies a maximal excursion of the entropy production, which enables us to write:
$P\left(\mathrm{max}_{t\leq\tau}S_{\mathrm{tot}}(t)\geq s\right)\leq e^{-ps}\left<e^{pS_{\mathrm{tot}}(\tau)}\right$>. (34)
On the other hand, in the case $e^{-pS_{\mathrm{tot}}}$, it increases as the entropy production decreases, and hence the probability of the exponential reaching some maximal value is related to the probability of the entropy production reaching some minimum value. Equivalently to the case +p+p, we can write for this case:
$P\left(\mathrm{min}_{t\leq\tau}S_{\mathrm{tot}}(t)\leq-s\right)\leq e^{-ps}\left<e^{-pS_{\mathrm{tot}}(\tau)}\right$>. (35)
These results provide the basis to our strategy based on maximal excursions of entropy production in Sec. III.3. They can also be used to find explicit expressions for optimal upper and lower thresholds that are not reached with a certain confidence level [48].
Appendix B Bounds for the output mechanical work
@@
$\begin{gathered}\eta=\frac{\left<W^{\mathrm{m}}(t)\right>}{\left<W^{\mathrm{c}}(t)\right>}=\frac{W_{\mathrm{ext}}}{\Delta\mu}.\end{gathered}$ (38)
Note that the macroscopic efficiency is time-independent. We can then write $\left<W^{\mathrm{c}}(t)\right>=\eta^{-1}\left<W^{\mathrm{m}}(t)\right$>, and express the chemical work in a trajectory up to a time $t$ by:
@@
being $\xi(t)$ a non-extensive in time stochastic variable with zero mean, $\left<\xi(t)\right>=0$. Going back to the total entropy production, using Eq. (39), we can now rewrite it as:
@@
[…]
@@
and where we denoted $W^{\rm m}_{\mathrm{max}}(t)=\max_{t\leq\tau}W^{\rm m}(t)$ the maximum of $W(t)$ within the interval $0,\tau]$. Equivalently, for the lower bound we have:
@@
and where we analogously denoted $W_{\min}^{\mathrm{m}}(t)=\mathrm{min}_{t\leq\tau}W^{\mathrm{m}}(t)$ the minimum of the mechanical work during the interval $0,\tau]$.
@@
which, after some basic algebra, give the following result for the optimal famility of thresholds not surpassed during the interval $0,\tau]$ with probability $1-\alpha$:
@@
Notice that the optimal upper bound still needs a reconstruction of the mean of the exponential of the entropy production through $\lambda(1)$. However, the optimal lower bound is achieved for p=1p=-1, and we have $\lambda(-1)=1$ from the standard IFT [89]. Therefore, in this case the bound saturates to a time-independent value:
@@
$\eta_{\rm syn}=\frac{\left<W^{\mathrm{c}}(t)\right>}{\left<W^{\mathrm{m}}(t)\right>}=\frac{\Delta\mu}{W_{\mathrm{ext}}}.$ (51)
Appendix C Experimental values for the parameters
Finally, the only available measurement of the time scale of the motor dynamics is its mean rotational speed. However, a full 360º rotation includes three net hydrolysis or synthesis steps in our model, which implies three cycles in either direction once. Given the lack of information about the rates of individual transitions, the first approximation was to consider $k_{\mathrm{ATP}}=k_{\mathrm{ADP}}=k_{HS}=\gamma$, with $\gamma$ the parameter that establishes the time scale of our simulations. Under normal conditions, the motor is expected to work in its stationary state, so we need the rate at which cycles are completed in the stationary regime and relate it to the experimental measurements. We obtained in Appendix D that the rate of cycles is given by Eq. (56). According to the first experimental characterization of the F1-ATPase in Ref. [18], in the absence of external forces ($W_{\mathrm{ext}}=0$) the motor rotates in the hydrolysis direction doing around $K_{\rm revs}=17$ r.p.s. We hence approximate the value of the rate $\gamma$ as:
$K_{x_{1}x_{2}}\pi_{x_{1}}-K_{x_{2}x_{1}}\pi_{x_{2}}=$
Appendix D Master equation and steady state